The annual salary for elementary teachers is left-skewed with a mean of $29,800 and a standard deviation of $2400. Suppose a sample of 100 teachers was chosen Describe the sampling distribution of the sample mean. O A. Normally distributed-s-$29,800, and ơřs24 OB. Approximately normally distributed, u -$29,800, and o $24 C. Normally distributed, --$29,800, and ơ -$240 0 D. Approximately normally distributed, μΧ-$29,800, and ơū_$240
Answer #1

Solution :

Given that,

mean = \mu = $29800

standard deviation = \sigma = $2400

n = 100

The sampling distribution of mean and standard deviation is ,

\mu\bar x = $29800 and

\sigma\bar x = \sigma / \sqrt n = 2400 / \sqrt 100 = 2400 / 101 = 240

Option D) is correct .